# Unveiling AI’s Blind Spots: An Oracle for In-Domain, Out-of-Domain, and Adversarial Errors

### v1

@@ 3.3 Datasets for training and testing mentors
Evaluation Metric.
To assess the performance of mentors, we report their error prediction accuracy on the test set corresponding to each specified error source. For instance, a mentor trained on the C10-ID training set is evaluated on the C10-OOD-SpN test set. The error prediction accuracy is calculated by averaging the mentor’s accuracies on the samples that the mentee correctly classified and those that the mentee incorrectly classified. However, since a mentee can make mistakes across various real-world scenarios, a mentor must accurately predict errors across all error types. Therefore, we compute the average accuracy of a mentor across all test sets, including one ID error, four OOD errors, and four AA errors. For simplicity, we refer to this average accuracy across all nine error sources as Accuracy. A mentor randomly guessing whether a mentee’s image classification is correct or incorrect for a given image would achieve an accuracy of 50%.
4 Results
A mentee’s mistakes can reveal their learning tendencies, behaviours, or traits. Here, we investigate which types of errors offer the most insight into understanding a mentee’s decision boundaries during image recognition tasks. We train mentors with identical architectures on datasets containing specific error types made by the mentee across C10 (Fig. 3(a)), C100 (Fig. 3(b)), and IN (Fig. 3(c)). For instance, if a mentor trained on C10-OOD achieves higher accuracy in error prediction compared to one trained on C10-IN, this suggests that in-domain errors provide less diagnostic information about the mentee’s decision-making process than out-of-domain errors. Both mentors and mentees may have the same or different backbones, such as ResNet50 (R) or ViT (V).
As shown in Fig. 3, over C10, C100, and IN images, the high accuracy for mentors trained on adversarial attack (AA) errors indicates that these errors offer deeper insights into the mentee’s decision process compared to out-of-domain (OOD) and in-domain (ID) errors. In some cases, mentors trained on OOD errors slightly outperformed those trained on ID errors, though both were still inferior to those trained on AA errors.
Loss landscape analysis.
A loss landscape of a mentee reflects how a mentee’s loss function behaves across different parameter configurations. Mentors’ performance offers insights into the structure of a mentee’s loss landscape. Consistent with (Ilyas et al., 2019), the high accuracy of mentors trained on AA errors suggests that adversarial images lie closer to the mentee’s decision boundary, enabling more accurate prediction of the mentee’s mistakes and a deeper understanding of the loss landscape. Similarly, OOD data aids mentors in learning decision boundaries by shifting ID samples closer to the boundary. However, it does not explore the boundary as thoroughly as adversarial images. ID data, with fewer samples near the boundary, provides more limited exploration compared to adversarial examples.
4.2 Mentor Architectures Matter in Error Predictions
Loss landscape analysis.
The performance difference between mentors’ architectures is due to ViT’s superior ability to identify features from error patterns. Its self-attention mechanism captures complex relationships among data samples, providing a deeper understanding of the mentee’s loss landscape, particularly in modelling irregular, rugged landscapes with sharp peaks and valleys.
Although adversarial images have been demonstrated to aid in error prediction (Sec. 4.1), it remains unclear whether adversarial images with varying degrees of image distortion exhibit the same effect. A straightforward method to regulate the level of image distortion caused by adversarial attacks is to set the perturbation bound . We employ four corruption levels by setting , , , and . We use the adversarial attack PIFGSM as an example since the error patterns from PIFGSM are most effective for the mentor’s prediction (see Fig. S1- S3). As shown in Fig. 5, the mentor’s accuracy significantly decreases as the distortion level increases. In particular, for the C10-AA-PIFGSM, the accuracy at level 1 is , which is notably higher than at level 4. Our findings suggest that adversarial attacks employing smaller perturbations yield more benefits for mentor error prediction. This phenomenon can be attributed to the fact that adversarial images with minimal perturbations maintain closer proximity to the decision boundary of a mentee.
4.4 Mentors generalize across mentees
In Sec. 4.1, mentors have demonstrated their ability to learn the error patterns of mentees. This observation raises an important question: can the error patterns learned from one mentee (mentee A) be generalized to another mentee (mentee B) when the two mentees employ different model architectures? To explore this, we evaluate all 324 mentors, whose performances are depicted in Fig. S1- S3, on the alternate mentee. Specifically, mentors trained on the errors of the ResNet50 mentee are tested on the predictions of the ViT mentee, and vice versa. The outcomes of these evaluations are illustrated in Fig. 5. Surprisingly, most points lie near the dashed diagonal line, implying that the mentors’ performance does not significantly deteriorate when evaluated on the predictions of different mentee architectures. This finding indicates that ResNet50 and ViT mentees tend to produce similar error patterns when trained on the same dataset.
4.5 Our Proposed SuperMentor Model outperforms other baselines
@@
| ✗ | ✗ | 57.5 | 61.0 | 56.1 | 58.6 | 54.3 | 58.6 | 59.1 | 58.5 | 59.6 | 58.2 | |
| ✗ | ✓ | 80.0 | 73.7 | 79.2 | 77.9 | 74.3 | 80.5 | 76.5 | 79.7 | 71.2 | 77.0 | |
| C10 | ours | 80.9 | 73.2 | 80.5 | 79.4 | 75.6 | 81.4 | 78.2 | 80.7 | 71.9 | 78.0 | |
| ✗ | ✗ | 56.8 | 59.5 | 56.6 | 57.8 | 53.7 | 57.7 | 57.3 | 57.3 | 57.1 | 57.1 | |
| ✗ | ✓ | 75.0 | 70.9 | 74.8 | 74.1 | 68.1 | 78.1 | 75.2 | 76.2 | 66.5 | 73.2 | |
| C100 | ours | 75.4 | 71.1 | 75.4 | 74.5 | 68.4 | 78.3 | 75.6 | 76.6 | 66.9 | 73.6 | |
| ✗ | ✗ | 73.0 | 70.1 | 69.6 | 72.8 | 68.5 | 75.8 | 72.5 | 73.6 | 70.7 | 71.9 | |
| ✗ | ✓ | 78.7 | 73.1 | 73.6 | 78.0 | 73.2 | 83.0 | 78.4 | 79.9 | 72.2 | 76.7 | |
| IN | ours | 78.9 | 73.6 | 74.6 | 78.3 | 73.6 | 83.0 | 78.4 | 79.9 | 72.3 | 77.0 | |
By drawing insights from observations in the subsections above, we propose an "oracle" mentor model, dubbed SuperMentor. We introduce the technical novelties of our SuperMentor below. First, as demonstrated in Sec. 4.1 and Sec. 4.3, mentors trained on adversarial images with small perturbations of a mentee outperform those trained on OOD and ID images; thus, our SuperMentor adopts the training data from the PIFGSM error source of mentees with . Second, since ViT has been proven to be a more effective architecture for mentors than ResNet50 (Sec. 4.2), SuperMentor adopts ViT as the backbone architecture.
Fig. 6 shows that SuperMentor outperforms other baseline mentors in the CIFAR-10 dataset. The detailed performance of the SuperMentor, along with other baseline mentors on various error sources from the CIFAR-10, CIFAR-100 and ImageNet-1K datasets, is presented in Fig. S1- S3.
@@
We also present the visualization of the SuperMentor’s embeddings on three types of error sources of a mentee in Fig. 7. It is evident that SuperMentor can effectively segregate samples correctly classified by the mentee from those that are misclassified, forming two distinct clusters.
Next, we examine the effect of the distillation loss (Fig. 2) on the SuperMentor performance. The results are presented in Tab. 2. It is clear that excluding results in a decrease in SuperMentor’s accuracy across all datasets. For example, in the C10 dataset, the average accuracy of SuperMentor decreases from to . This suggests that encourages SuperMentor to learn the fine-grained decision boundaries among different object classes of a mentee.
Alternatively, instead of utilizing the mentee’s logits, SuperMentor can incorporate an additional cross-entropy loss to align the mentor’s predicted object class labels with those of the mentee, denoted as . From Tab. 2, we observe that replacing with leads to a slight decrease in accuracy. This is due to the fact that the mentee’s logits contain more information than the mentee’s class labels.
5 Conclusion
In our work, we tackle the challenge of predicting errors in AI models through extensive empirical evaluations using an end-to-end trainable "mentor" model. This mentor model is designed to assess the correctness of a mentee model’s predictions across three distinct error types: in-domain errors, out-of-domain errors, and adversarial attack errors. Our results show that the mentor model excels at learning from a mentee’s errors on adversarial images with minimal perturbations and, surprisingly, generalizes well to both in-domain and out-of-domain predictions of the same mentee. Additionally, we highlight the effectiveness of transformer-based mentor architectures compared to 2D-CNN-based ones, demonstrating their superior generalization capabilities across mentees with diverse backbones. Lastly, we introduce the SuperMentor, which significantly outperforms all existing mentor baselines.
Our work paves the way for several promising research directions in the field of safe and trustworthy AI. First, while our current research focuses on image classification, there is potential to extend this approach to other vision and language tasks, such as object detection and machine translation. Second, future research could explore mutual learning between mentors and mentees, where mentors not only learn from the mentee’s error patterns but also provide valuable feedback to help refine the mentee. Third, we can establish more rigorous evaluation criteria for mentors, broadening their predictive capabilities. For example, beyond predicting whether a mentee is likely to make errors, mentors could also forecast the specific types of errors a mentee may encounter. Fourth, this concept can be applied to investigate recognition errors in humans and primates, drawing parallels with AI models. Such analysis could provide insights into error pattern alignment between biological and artificial intelligent systems. Overall, our work lays the foundation for developing systems capable of anticipating the errors of others, offering practical value in high-stakes real-world applications.

### v2

@@ Appendix A Dataset split for each error source used in mentor training
Appendix B Details of baselines
For performance comparison, we adopt seven baselines, as introduced in Sec. 3.4. Their details are outlined below.
Self Error Rate (SER): We predict the correctness of the mentee’s outputs by referencing its accuracy on in-domain samples. For instance, if the mentee’s in-domain accuracy is 70%, we use a random binary generator that produces 1 (“correct”) with 70% probability and 0 (“wrong”) with 30% probability, applying this to all testing scenarios.
Maximum Class Probability (MCP) (Hendrycks & Gimpel, 2016): As mentioned in Sec. 3.1, the mentee’s output logit is denoted as . The maximum class probability of the mentee’s output is defined as where is the softmax function. The MCP indicates how confident the mentee is in its most probable prediction. We set a threshold to distinguish between the mentee’s confident and unconfident predictions. Specifically, if , the mentee’s prediction will be considered a correct prediction due to its high confidence; otherwise, it will be regarded as an incorrect prediction. We employ three predefined thresholds for : 0.5, 0.7, and 0.9.
Class Probability Entropy (CPE): The class probability entropy of the mentee’s output is defined as , where denotes the softmax function and represents the entropy measure quantifying the uncertainty in the probability distribution. A high entropy value signifies a high level of uncertainty in the mentee’s predictions. The entropy reaches its maximum value, , when the mentee’s class probabilities in are equal. We define the uncertainty threshold as , where . If , the mentee’s prediction is considered correct, indicating sufficient certainty in its prediction. Otherwise, the prediction is regarded as incorrect. We set three predefined values for : 0.01, 0.1 and 0.3.
Distance to centroid (DTC): The embedding generated by the mentee before the final binary classification layer represents the mentee’s feature interpretation of each sample. To determine the feature centroid for each class, we first average the features of all testing samples within that class based on the mentee’s predictions. Next, we calculate the L2 distance between each sample’s feature and its corresponding class centroid, denoted as . We establish a distance threshold ; if , the mentee’s prediction is considered correct since the sample is close to the class centroid. Otherwise, the prediction is deemed incorrect. We have set three predefined values for : 10, 20, and 30.
ConfidNet (Corbière et al., 2019): A shallow failure prediction neural network that learns True Class Probability (TCP) from the training set instead of relying on Maximum Class Probability (MCP).
TrustScore (Jiang et al., 2018): It introduces a new score called trust score by measuring the agreement between the classifier and a modified nearest-neighbor classifier on the testing samples for error prediction.
Steep Slope Loss (SSL) (Luo et al., 2021): Training an AI model for trustworthiness prediction by leveraging a carefully-designed steep slope loss function.
Appendix C Example of calculating a mentor’s average accuracy using the proposed evaluation metric
As an example of computing the accuracy of a mentor evaluated using the metric presented in Sec. 3.4, we consider a mentor trained on C10-AA-PIFGSM and evaluate it on all the following datasets with their accuracies of 10% on C10-ID, 20% on C10-OOD-SpN, 30% on C10-OOD-GaB, 40% on C10-OOD-Spat, 50% on C10-OOD-Sat, 60% on C10-AA-Jitter, 70% on C10-AA-PGD, 80% on C10-AA-CW, and 90% on C10-AA-PIFGSM testing samples (See Appendix, Tab. S1 for testing splits), the average accuracy of this mentor is calculated as (10% + 20% + 30% + 40% + 50% + 60% + 70% + 80% + 90%) / 9 = 50%. For simplicity, we refer to this average accuracy across all nine error sources as Accuracy. A mentor randomly guessing whether a mentee’s image classification is correct or incorrect for a given image would achieve an accuracy of 50%.
Appendix D Joint training
As mentioned in Sec. 4.1, we add an experiment by training a Vision Transformer (ViT)-based mentor model on the correctness of a mixture of ID, OOD, and AA data for the ResNet-based mentee model from the CIFAR-10 dataset. The OOD data was corrupted with speckle noise, and the AA data was generated using PIFGSM. Compared to training the mentor solely on AA data which achieved an average accuracy of 78.0%, including ID and OOD data improved the average accuracy slightly to 79.1%. This minor improvement suggests a significant overlap of mentee error patterns among ID, OOD, and AA data. In other words, training mentors on OOD or ID data do not necessarily benefit mentors. Therefore, considering the associated computational costs, it is recommended to train mentors exclusively on AA data in practical applications. This finding underscores the significance of our study by providing critical insights into the optimal training practices for mentor models.
Appendix E More comparisons between SuperMentor and baselines
As noted in the caption of Tab. 1, baseline performances across multiple hyperparameter configurations are indicated in Appendix, Tab. S2. In addition, when ViT serves as the mentee, the SuperMentor’s performance is shown in Appendix, Tab. S3. It can be seen that our SuperMentor can still achieve the best performance when the mentee has a backbone of ViT. This emphasizes the importance of the error source used during training.
| ID | SpN | GaB | Spat | Sat | PGD | CW | Jitter | PIFGSM | Average | |
| SER | 50.2 | 49.8 | 49.9 | 50.1 | 49.9 | 50.1 | 49.9 | 49.9 | 49.9 | 50.0 |
| MCP () | 69.5 | 73.5 | 73.6 | 72.2 | 71.7 | 66.0 | 66.7 | 65.1 | 64.9 | 69.2 |
| MCP () | 77.3 | 76.8 | 77.4 | 78.3 | 76.6 | 71.8 | 72.9 | 71.7 | 69.4 | 74.7 |
| MCP () | 78.2 | 73.3 | 74.5 | 77.0 | 74.9 | 72.2 | 72.3 | 73.0 | 69.0 | 73.8 |
| CPE () | 69.1 | 62.1 | 63.4 | 66.0 | 64.0 | 63.7 | 62.6 | 65.1 | 60.7 | 64.1 |
| CPE () | 78.0 | 75.2 | 75.9 | 78.0 | 76.2 | 72.2 | 72.8 | 72.7 | 69.3 | 74.5 |
| CPE () | 64.0 | 70.1 | 69.6 | 67.6 | 67.0 | 61.2 | 61.5 | 60.6 | 60.8 | 64.7 |
| DTC () | 52.4 | 51.2 | 52.9 | 52.7 | 52.5 | 51.4 | 51.6 | 51.7 | 51.6 | 52.0 |
| DTC () | 52.8 | 51.1 | 50.5 | 51.4 | 51.2 | 52.8 | 51.9 | 52.0 | 52.7 | 51.8 |
| DTC () | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 |
| ConfidNet (Corbière et al., 2019) | 67.3 | 65.9 | 68.5 | 70.8 | 67.9 | 68.8 | 69.0 | 70.1 | 67.1 | 68.4 |
| TrustScore (Jiang et al., 2018) | 70.0 | 67.5 | 71.1 | 73.4 | 71.3 | 73.2 | 72.5 | 73.9 | 72.7 | 71.7 |
| SSL (Luo et al., 2021) | 66.3 | 66.7 | 69.7 | 71.9 | 70.0 | 69.1 | 70.1 | 70.2 | 66.7 | 69.0 |
| SuperMentor (ours) | 78.9 | 73.6 | 74.6 | 78.3 | 73.6 | 83.0 | 78.4 | 79.9 | 72.3 | 77.0 |
| ID | SpN | GaB | Spat | Sat | PGD | CW | Jitter | PIFGSM | Average | |
| SER | 49.9 | 50.0 | 49.9 | 49.9 | 50.0 | 49.8 | 50.0 | 49.9 | 50.0 | 49.9 |
| MCP () | 69.9 | 72.9 | 71.5 | 71.5 | 71.8 | 60.5 | 63.6 | 56.6 | 60.7 | 66.6 |
| MCP () | 78.5 | 77.9 | 78.0 | 78.1 | 78.0 | 63.9 | 67.8 | 61.5 | 63.8 | 71.9 |
| MCP () | 58.0 | 53.3 | 57.2 | 55.5 | 57.6 | 50.6 | 51.6 | 50.9 | 49.2 | 53.8 |
| CPE () | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 |
| CPE () | 51.4 | 50.7 | 51.2 | 51.1 | 51.5 | 49.6 | 50.1 | 49.6 | 49.1 | 50.5 |
| CPE () | 72.4 | 73.7 | 73.5 | 72.7 | 73.5 | 61.3 | 63.7 | 57.0 | 60.2 | 67.5 |
| DTC () | 69.5 | 68.9 | 69.3 | 68.6 | 68.6 | 65.7 | 64.1 | 61.8 | 67.0 | 67.1 |
| DTC () | 50.9 | 50.4 | 50.4 | 50.6 | 50.6 | 50.2 | 50.3 | 50.2 | 50.3 | 50.4 |
| DTC () | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 |
| ConfidNet (Corbière et al., 2019) | 73.3 | 74.1 | 73.4 | 75.6 | 74.5 | 69.2 | 70.5 | 66.9 | 70.5 | 72.0 |
| TrustScore (Jiang et al., 2018) | 73.7 | 72.0 | 72.4 | 73.5 | 72.5 | 70.4 | 70.1 | 69.2 | 72.5 | 71.8 |
| SSL (Luo et al., 2021) | 64.3 | 66.3 | 68.2 | 70.6 | 69.3 | 62.5 | 66.0 | 60.1 | 62.4 | 65.5 |
| SuperMentor (ours) | 72.7 | 71.3 | 71.7 | 73.0 | 70.2 | 80.2 | 74.1 | 72.9 | 74.1 | 73.4 |
Appendix F Ablation study
As mentioned in Sec. 4.5, we examine the effect of the distillation loss (Fig. 2) on the SuperMentor performance. The results are presented in Appendix, Tab. S4. It is clear that excluding results in a decrease in SuperMentor’s accuracy across all datasets. For example, in the C10 dataset, the average accuracy of SuperMentor decreases from to . This suggests that encourages SuperMentor to learn the fine-grained decision boundaries among different object classes of a mentee.
Alternatively, instead of utilizing the mentee’s logits, SuperMentor can incorporate an additional cross-entropy loss to align the mentor’s predicted object class labels with those of the mentee, denoted as . From Appendix, Tab. S4, we observe that replacing with leads to a slight decrease in accuracy. This is due to the fact that the mentee’s logits contain more information than the mentee’s class labels.
@@ Appendix F Ablation study
Appendix G Visualization
@@
To gain an intuitive understanding of SuperMentor’s binary classification performance in error prediction, as mentioned in Sec. 4.5, we present the visualization of SuperMentor’s embeddings on three types of error sources of a mentee in Appendix, Fig. S1. It is evident that SuperMentor can effectively segregate samples correctly classified by the mentee from those that are misclassified, forming two distinct clusters.
Appendix H SuperMentor in real-world practice
To reflect the real-world contribution of our work, as mentioned in Sec. 5, we expanded our experiments to the NCTCRCHE100K (Kather et al., 2019) dataset which is used for medical image classification on colorectal cancer. This dataset comprises around 100K images across 9 tissue classes. Predicting AI models’ errors in medical image classification has significant real-world implications, such as reducing misdiagnoses and increasing the reliability of AI-assisted medical tools. Tab. S5 demonstrates that our SuperMentor accurately predicts the correctness of mentee outputs with an average accuracy of 81.9%, significantly outperforming other baseline methods. This highlights that our proposed framework can offer greater value and reliability for AI error prediction compared to existing approaches in medical image classification.
Given the increasing integration of AI models into our daily lives, ensuring their accuracy is paramount, especially in high-stakes fields such as medicine and finance. This experiment with medical image datasets underscores the critical role that SuperMentor plays in enhancing reliability and trustworthiness in these vital areas, showcasing its potential to make a meaningful impact where precision and dependability are essential.
| ID | SpN | GaB | Spat | Sat | PGD | CW | Jitter | PIFGSM | Average | |
| SER | 49.8 | 49.6 | 50.4 | 50.0 | 49.9 | 50.1 | 50.2 | 49.8 | 49.8 | 50.0 |
| MCP () | 53.9 | 59.4 | 54.3 | 54.6 | 50.4 | 54.2 | 54.4 | 54.7 | 55.0 | 54.6 |
| MCP () | 65.1 | 71.3 | 65.0 | 66.2 | 49.9 | 66.2 | 66.1 | 65.8 | 67.1 | 64.7 |
| MCP () | 74.5 | 78.7 | 74.3 | 76.2 | 49.4 | 76.9 | 75.8 | 75.9 | 76.2 | 73.1 |
| CPE () | 78.5 | 74.2 | 81.9 | 77.7 | 45.2 | 77.4 | 77.4 | 77.8 | 77.3 | 74.2 |
| CPE () | 77.6 | 79.3 | 77.8 | 78.2 | 48.8 | 79.2 | 78.0 | 78.8 | 78.3 | 75.1 |
| CPE () | 67.6 | 76.3 | 67.0 | 68.8 | 51.7 | 70.3 | 69.0 | 69.9 | 69.5 | 67.8 |
| DTC () | 77.7 | 67.3 | 77.2 | 79.7 | 66.2 | 78.8 | 77.4 | 77.6 | 78.1 | 75.5 |
| DTC () | 69.4 | 48.5 | 67.2 | 61.6 | 52.4 | 59.2 | 64.7 | 62.2 | 59.2 | 60.5 |
| DTC () | 52.6 | 49.4 | 53.2 | 50.7 | 50.0 | 51.1 | 52.4 | 51.5 | 50.9 | 51.3 |
| ConfidNet (Corbière et al., 2019) | 83.6 | 59.7 | 89.4 | 74.0 | 56.5 | 80.1 | 85.7 | 82.1 | 80.4 | 76.8 |
| TrustScore (Jiang et al., 2018) | 73.3 | 62.9 | 75.2 | 63.7 | 63.7 | 73.1 | 74.9 | 73.5 | 74.0 | 70.5 |
| SSL (Luo et al., 2021) | 79.3 | 57.1 | 87.4 | 75.9 | 56.9 | 76.2 | 83.1 | 77.7 | 76.9 | 74.5 |
| SuperMentor (ours) | 88.7 | 64.9 | 69.8 | 85.0 | 54.5 | 89.4 | 88.7 | 88.2 | 87.4 | 81.9 |
Appendix I SuperMentor performance on additional OOD domains on ImageNet dataset
As mentioned in Sec. 5, we included experiments to evaluate the SuperMentor in more diversified OOD domains on ImageNet datasets, including ImageNet9-MIXED-RAND (IN9-MR) (Xiao et al., 2020), ImageNet9-MIXED-SAME (IN9-MS) (Xiao et al., 2020), ImageNet9-MIXED-NEXT (IN9-MN) (Xiao et al., 2020), ImageNet-R (IN-R) (Hendrycks et al., 2021) and ImageNet-Sketch (IN-S) (Wang et al., 2019). Specifically, the MIXED-RAND, MIXED-SAME, and MIXED-NEXT datasets are derived from 9 classes in ImageNet and contain varying amounts of background and foreground signals. These datasets aim to demonstrate that models often classify objects based on background cues (often spurious features), rather than the objects themselves. Specifically, MIXED-SAME, MIXED-RAND, and MIXED-NEXT represent images with random backgrounds from the same class, random backgrounds from a random class, and random backgrounds from the next class, respectively. The ImageNet-R dataset comprises images featuring artistic representations of objects, such as cartoons, community-generated art, and graffiti renditions. The ImageNet-Sketch dataset consists of sketch-like images that match the ImageNet validation set in both categories and scale.
To achieve a more thorough evaluation of our SuperMentor (see Sec. 4.5) in OOD domains, we assess its performance on these additional OOD datasets. It is worth noting that the SuperMentor learns from the mentee’s errors on adversarial ImageNet images generated by the PIFGSM attack only. In other words, the SuperMentor was NOT trained on any of these additional ImageNet-related OOD datasets. The SuperMentor achieves average accuracies of 68.46%, 69.24%, 67.04%, 58.68%, and 58.68% on the IN9-MR, IN9-MS, IN9-MN, IN-R, and IN-S datasets respectively, surpassing the 50% chance level. Despite that our Supermentor model is a simple ViT, it is still quite remarkable to achieve above-chance error prediction performance. We hope our work inspires researchers to explore more sophisticated AI mentors, incorporating advanced architectures or loss functions, to further enhance the ability to predict the correctness of another AI model’s outputs.
Appendix J Discussion on the vulnerabilities of the mentor AI
Beyond error prediction, our framework improves the understanding of both individual and systemic vulnerabilities in Deep Neural Network (DNN)-based mentees, fostering more robust AI systems. This is further illustrated by the EigenCAM (Gildenblat & contributors, 2021; Muhammad & Yeasin, 2020) visualization in Appendix, Fig. S2. From these visualizations, we can make two observations below:
The vulnerabilities of the mentee do not overlap with those of the mentors. As illustrated in Appendix, Fig. S2, the mentor does not merely replicate the mentee’s learning patterns, as their activation maps for identical input images can differ significantly. For example, in the images shown in Appendix, Fig. S2 (a), although the mentor and mentee exhibit very similar activation maps for ID images, their activations diverge greatly when the input image is corrupted by Gaussian Blur (GaB) or subjected to adversarial attacks using PIFGSM method. In the IN-OOD-GaB and IN-AA-PIFGSM images, the mentee demonstrates vulnerability under both corruption and adversarial attacks, whereas our SuperMentor does not exhibit such vulnerabilities on these OOD and AA images, consistently focusing on the brambling object. This indicates the vulnerabilities of mentors and mentees are distinct due to their non-overlapping objectives.
Our AI mentor serves as a valuable diagnostic tool for AI mentees. We argue that, beyond merely detecting errors, mentors function as important diagnostic tools for AI mentees, as supported in Appendix, Fig. S2. For example, in Appendix, Fig. S2 (b), although the mentee correctly classifies the ID image, its activation map does not concentrate on the baseball but instead highlights background cues (spurious features). When the input image is corrupted by Gaussian Blur (GaB) or subjected to adversarial PIFGSM attacks, the mentee fails to make accurate predictions. In contrast, our SuperMentor effectively focuses on the stitches of the baseball—a key feature—regardless of whether the image is corrupted or adversarially attacked. The mentor highlights the stitches of the baseball as the key feature pattern that the mentee tends to overlook. This suggests that the mentor can identify vulnerabilities in the DNN-based model, even though it shares the same ViT architecture as the mentee. Furthermore, this reveals significant potential for our mentor to help refine the mentee’s performance by providing valuable insights into the mentee’s behavior.
In conclusion, Appendix, Fig. S2 provides strong evidence that the AI mentor not only does not share the same vulnerabilities as the mentee but also serves as a valuable diagnostic tool for AI mentees.